Professor plays basketball and makes 75% of free throws she shoots. If professor shot 5 free throws today, what is the probability that she makes at least 1?
Answer #1

Solution

Given that ,

p = 0.75

q = 1 - p = 1 - 0.75 = 0.25

n = 5

Using binomial probability formula ,

P(X = x) = (n C x) * p x * (1 - p)n - x

P(X 1  ) = 1 - P( x <1)

= 1 - P(X = 0)

= 1 - (5 C 0) * 0.75 0 * (0.25)5

=1-0.0010

probability=0.9990